# Logic Jailbreak: Efficiently Unlocking LLM Safety Restrictions Through Formal Logical Expression

### v1

@@ 3.1 Dataset Construction
3.2 Evaluation
To evaluate the effectiveness of our jailbreak method, we measure the Attack Success Rate (ASR), where ASR@N denotes the success rate over N attack attempts. To provide a comprehensive assessment of Logibreak’s performance, we adopt three complementary evaluation methods, following the approach of Andriushchenko and Flammarion (2024).
The first method is a rule-based evaluation proposed by Zou et al. (2023), which identifies specific words or phrases that indicate the target language model has refused to respond. While useful, this approach is limited, as it does not consider the semantic content of the response—an essential aspect in determining whether the output is potentially harmful. To address this shortcoming, we additionally adopt the semantic evaluation protocols from Chao et al. (2023) and Chao et al. , using GPT-4 and Llama-3 70B to assess outputs based on the same prompts employed in those studies.
For the Chinese and Dutch settings, we translated the corresponding prompts using GPT-3.5, with all translations manually reviewed to ensure semantic accuracy and contextual appropriateness.
3.3 Baselines
To evaluate the effectiveness of our proposed approach, we compare it against six SOTA black-box jailbreak methods from the literature. Detailed descriptions are provided in Appendix C.
3.4 Models
We evaluate our method across a diverse set of target models, encompassing both open-source and closed-source LLMs. The open-source models include Qwen-2.5-7B Yang et al. (2024), Llama-3-8B Grattafiori et al. (2024), DeepSeek V3 Liu et al. (2024), and DeepSeek R1 Guo et al. (2025), while the closed-source models consist of GPT-3.5 OpenAI (2023) and GPT-4o-mini OpenAI (2024). For all models, we adopt a black-box setting, wherein we interact with the models exclusively through their APIs.
This setup ensures a fair comparison across different models and allows us to evaluate the generalizability of LogicBreak across diverse architectures under the most stringent conditions.
Table 3: LogiBreak jailbreak performance on Chinese and Dutch.
Chinese Dutch
Rule_Judge Llama3-70B_Judge GPT4-Judge Rule_Judge Llama3-70B_Judge GPT4-Judge
ASR@1 ASR@5 ASR@1 ASR@5 ASR@1 ASR@5 ASR@1 ASR@5 ASR@1 ASR@5 ASR@1 ASR@5
Llama 100 100 42 48 56 69 100 100 89 89 87 93
Qwen 90 94 31 48 67 88 92 96 86 90 90 96
DS-V3 87 99 63 80 79 93 96 100 90 95 89 98
GPT-3.5 92 100 58 81 80 96 97 100 79 88 90 98
GPT-4o 90 100 50 75 75 93 93 100 80 91 88 95
Avg. 91.8 98.6 48.8 66.4 71.4 87.8 95.6 99.2 84.8 90.6 88.8 96.0
3.5 Overall Results
The empirical results in Table 1 show that LogiBreak consistently improves ASR across various evaluation frameworks and judging models. Specifically, it achieves average ASR@5 gains of over 4.9×, 35.3×, and 17.8× under rule-based, Llama, and GPT judges, respectively, demonstrating the robustness and generalizability of logically guided transformations in bypassing safety constraints. Notably, even under the stricter ASR@1 setting—where only the first adversarial attempt counts—LogiBreak maintains an average ASR exceeding 40% across all judges, suggesting that a single transformation often suffices. Compared to baselines, LogiBreak outperforms all other methods across all target LLMs and metrics, as shown in Table 2. It is noteworthy that although Caesar achieved a 100% ASR under the rule-based judge, its ASR was significantly lower under other semantic judges. This indicates that the LLM failed to accurately perform the encoding and decoding tasks, resulting in outputs that lack meaningful content.
@@ 3.5 Overall Results
3.6 Result on Different Languages
The results of LogiBreak on Chinese and Dutch inputs are presented in Table 3. On Chinese inputs, we observe that although the attack success rate under the rule-based judge is relatively high—reaching 90%—there is a notable performance degradation under semantic-based judges. Specifically, models such as Llama3-8B and Qwen-2.5-7B exhibit a significant drop in ASR@5, falling below 50% when evaluated by the Llama judge. Upon further inspection of failure cases, we find that smaller LLMs occasionally struggle to comprehend logical expressions in Chinese. In such cases, although the model does not explicitly refuse to respond, its answers often fail to fulfill the actual intent of the request, indicating a mismatch in semantic understanding.
In contrast, LogiBreak achieves the highest ASR@5 on Dutch inputs across all judges, even reaching 90% under each evaluation setting. We attribute this strong performance to the limited safety alignment resources available for low-resource languages like Dutch. This finding is consistent with previous work Yong et al. (2023), which suggests that languages with limited alignment data tend to exhibit weaker safety.
@@ 3.7 Failure Analysis
By analyzing the attack success rates across categories in Figure 4, we observe distinct patterns of vulnerability when applying LogiBreak to various LLMs. Specifically, the Fraud/Deception and Privacy categories consistently exhibit high penetration rates—exceeding 80% across all models and evaluation setups—indicating that these areas are particularly susceptible to logical adversarial prompts. In contrast, the Sexual/Adult Content and Expert Advice categories demonstrate greater resilience, with success rates reliably falling below the 80% threshold.
@@
Furthermore, the variation in attack success rates appears correlated with the perceived severity of harm associated with each category. Categories that have attracted greater public concern—such as explicit content or risky advice—tend to show stronger alignment, likely due to their more prominent representation in alignment datasets. This trend aligns with findings from Chen et al. (2024) and points to an imbalanced safety landscape in current LLMs, where vulnerabilities are not uniformly addressed across harm domains.
@@ 3.8 Ablation Study
In our ablation study, we evaluated logical transformation’s role in bypassing safety alignment in LLMs. When the natural-language request accompanied its logical form, ASR@5 dropped substantially versus using the logical form alone (Figure 5), confirming that logical expressions lie outside the model’s alignment distribution ($\mathcal{X}_{\text{logic}}\cap\mathcal{D}_{\text{aligned}}\approx\emptyset$). Models detect and curb harmful patterns in natural language but fail to do so when intent is encoded purely as logic.
We also assessed the prepended contextual grounding phrase $x_{context}$. Removing it consistently raised ASR across evaluations (Figure 5), showing that this phrase crucially frames the logical expression as valid input rather than a safety violation. Together, these findings validate LogiBreak’s mechanism and expose alignment vulnerabilities to semantic-preserving transformations.
Figure 5: Ablation study of the prepended contextual grounding phrase and the role of logical transformation by introducing natural-language requests.
@@ 3.9 Mitigation
Table 4: Performance of prompt-based Self-Reminder (S_R) and finetune-based Llama-Guard (Guard) mitigations against LogiBreak.
Rule_Judge Llama3-70B_Judge
LogiBreak S_R Guard LogiBreak S_R Guard
Llama 82 17 32 70 6 22
Qwen 78 81 33 81 70 28
DS-V3 88 90 35 84 40 25
GPT-3.5 82 76 35 83 49 26
GPT-4o 82 74 34 85 54 29
We evaluate LogiBreak against two mitigation strategies: the prompt-based Self-Reminder Xie et al. (2023) technique and the fine-tuned model Llama-Guard-3-8B Llama Team (2024) as filter.
For Self-Reminder mitigation against LogiBreak, effectiveness varies significantly across models. Under Rule_Judge, only Llama3-8B shows substantial improvement (ASR drop: 82% to 17%), while other models (Qwen-2.5, DeepSeek-V3, GPT-3.5, GPT-4o-mini) remain vulnerable. With Llama3-70B_Judge, Llama3-8B achieves the greatest improvement (70% to 6%), while others show modest gains—Qwen remains particularly vulnerable (70% ASR), and DeepSeek-V3, GPT-3.5, and GPT-4o maintain ASRs of 40-54%.
For Llama-Guard defense, despite being fine-tuned specifically for content safety classification, approximately 36% of LogiBreak prompts successfully bypass the guard and are incorrectly classified as safe. These "safe" inputs still achieve attack success rates exceeding 30% under Rule_Judge and 20% under Llama_Judge evaluation across various models.
@@
3.10 t-SNE visualization
Figure 6: t-SNE projection of LLM2Vec embeddings computed for three inputs: the original natural-language request, its reformulation as a logical expression, and the whole input prompt.
Figure 6 presents a t-SNE Van der Maaten and Hinton (2008) projection of LLM2Vec BehnamGhader et al. embeddings computed for three inputs: the original natural-language request, its reformulation as a logical expression, and the whole input prompt. In both English and Chinese settings, these three embedding types form clearly separable clusters; however, the clustering is markedly sharper in the English case. We suggest that English translation systematically converts multi-word phrases into camel-cased atomic tokens (e.g., "social media campaign" to "SocialMediaCampaign"), creating semantically distinct lexical units that lack direct Chinese equivalents due to the logographic nature of Chinese script. This tokenization asymmetry manifests in the embedding space as tighter intra-cluster cohesion for English transformed tokens compared to their Chinese counterparts. These findings demonstrate that LogiBreak can successfully transform the embedding distribution, enabling effective jailbreak.
@@ 4.1 First order logic in NLP
First-order logic (FOL) plays a crucial role in many logic-based NLP applications, making the task of translating natural language to FOL (NL-FOL) particularly important. In tasks like Recognizing Textual Entailment (RTE), several works have leveraged FOL to model natural language entailment Bos and Markert (2005); Lee et al. (2025). Similarly, for natural logic inference (NLI), FOL has been employed to derive commonsense conclusions Angeli and Manning (2014). Moreover, in the context of knowledge graphs, FOL enables complex query answering through embedding-based neuralization of logical operators Xia et al. (2025); Guu et al. (2015); Hamilton et al. (2018).
Early efforts in NL-to-FOL translation primarily relied on grammar-based techniques Purdy (1991); Angeli and Manning (2014); MacCartney and Manning (2014) or neural networks Singh et al. (2020); Lu et al. (2022). With the rapid progress of LLMs, however, researchers have increasingly explored LLM-based solutions. For instance, LogicLLaMA, a fine-tuned Llama variant introduced in Xiong et al. (2024), is designed specifically for NL-to-FOL conversion. The NL2FOL framework uses LLMs to convert natural language into FOL, enabling tasks like logical fallacy detection through satisfiability checking Lalwani et al. (2024). Despite these advances, the safety of FOL for LLMs remains an unexplored area.
@@ 4.2 Jailbreak attack
As LLMs gain widespread use, they have been shown to be vulnerable to jailbreak attacks—adversarial prompts that induce harmful or restricted outputs. These attacks fall into two main types: white-box methods, which require access to model internals (e.g., logits, gradients) and are limited to open models; and black-box methods, which exploit only input–output behavior Yi et al. (2024). Due to the impracticality of white-box access, we focus on the black-box setting.
Within black-box settings, a major line of work involves prompt injection attacks, which embed adversarial instructions or tokens into user inputs to bypass alignment constraints Shen et al. (2024); Wei et al. (2023); Liu et al. (2023); Zhou et al. (2024a); Yu et al. (2024). While often effective, these methods are highly sensitive to system prompt structures and safety mechanisms, making them susceptible to mitigation through prompt updates, pattern detection, or improved privilege classification. As a result, prompt injection has become a reactive, unstable strategy whose effectiveness is frequently undermined by model and infrastructure updates.
Another important class comprises black-box optimization methods, which frame adversarial prompting as a discrete optimization problem under query-only constraints. Techniques such as deep Q-learning Jawad and BRUNEL (2024) and latent Bayesian optimization Basani and Zhang (2024) search for token sequences or suffixes that elicit harmful outputs while avoiding moderation. However, these approaches often suffer from low efficiency, requiring large query budgets and extensive search to achieve success.
Recent work also explores language-based vulnerabilities, revealing that LLMs may exhibit reduced safety and robustness when interacting in certain low-resource languages Yong et al. (2023); Deng et al. . While promising, these findings remain limited by the small number of languages examined.
@@ 5 Conclusion
We propose LogiBreak, a novel black-box jailbreak method that transforms malicious natural language into formal logical expressions, preserving semantic intent while evading detection by alignment mechanisms. Through experiments on multiple LLMs across three languages, we demonstrate the effectiveness of our approach, achieving high success rates with minimal attempts. The success of LogiBreak highlights a critical vulnerability in current LLMs—namely, the lack of semantic-level safety alignment. Our findings emphasize the need for post-training alignment that operates at the semantic level to ensure more robust model safety.
6 Limitation
Our primary limitation lies in the constrained scope of our experimental evaluation, particularly the inability to extend testing to a broader range of jailbreak datasets or additional languages. Although Logibreak has demonstrated stability and effectiveness across various tasks, limited resources have restricted the scale and diversity of our assessments. A more comprehensive evaluation across diverse datasets and multilingual scenarios would offer stronger evidence of Logibreak’s generalizability, which we leave as an important direction for future work.
Appendix A Implementation of the Judges
In Figures 8, 9, and 10, we showcase the prompts or functions used for the Rule-based, LLaMA, and GPT judges in three different languages.
Appendix B LogiBreak Prompt
@@
Appendix C Baselines
  • Casaer is a cipher-based jailbreak technique proposed in Yuan et al. , which involves training the target LLM on the principles of encoding and decoding, and subsequently employing encoded malicious inputs to attempt to bypass its safety alignment mechanisms.
Figure 7: LogiBreak Prompt in English, Chinese and Dutch.
Appendix D Request Categories and Abbreviations in the Dataset
The request categories and their corresponding abbreviations in our dataset are presented in Table X, following the taxonomy proposed by Chao et al. .
Table 5: Categories and corresponding Abbreviations.

### v2

@@ 3.5 Overall Results
Table 3: Performance comparison with five black-box jailbreak baselines measured by ASR@5 (in %).
@@
Rule_Judge LLaMA_Judge GPT_Judge
ASR @1 @5 @1 @5 @1 @5
ZH Llama 100 100 42 48 56 69
Qwen 90 94 31 48 67 88
DS-V3 87 99 63 80 79 93
GPT-3.5 92 100 58 81 80 96
GPT-4o 90 100 50 75 75 93
Avg. 91.8 98.6 48.8 66.4 71.4 87.8
DU Llama 100 100 89 89 87 93
Qwen 92 96 86 90 90 96
DS-V3 96 100 90 95 89 98
GPT-3.5 97 100 79 88 90 98
GPT-4o 93 100 80 91 88 95
Avg. 95.6 99.2 84.8 90.6 88.8 96.0
JA Llama 100 100 80 91 22 39
Qwen 99 100 91 92 21 37
DS-V3 94 100 91 97 33 58
GPT-3.5 96 99 89 98 44 65
GPT-4o 97 100 90 97 42 66
Avg. 97.2 99.8 88.2 95.0 32.4 53.0
ES Llama 96 96 86 86 71 79
Qwen 92 93 94 94 68 80
DS-V3 94 99 96 98 78 81
GPT-3.5 92 99 89 95 60 88
GPT-4o 93 98 93 96 67 83
Avg. 93.4 97.0 91.6 93.8 68.8 82.2
Table 4: LogiBreak’s performance across four additional languages (in %).
3.6 Result on Additional Languages
The results of LogiBreak on Chinese, Dutch, Japanese and Spanish inputs are presented in Table 4. We observed a noteworthy performance drop for smaller LLMs like LLaMA3-8B and Qwen-2.5-7B on Chinese and Japanese inputs, particularly when evaluated by semantic-based judges like LLaMA Judge and GPT Judge. While the ASR reached 90% with a rule-based judge for Chinese, the ASR@5 plummeted below 50% under the LLaMA judge. Upon analysis, we found that these smaller LLMs struggle to comprehend complex logical expressions in both Chinese and Japanese. Even when they don’t explicitly refuse to respond, their answers often miss the request’s true intent, highlighting a semantic understanding mismatch.
In comparison, LogiBreak achieves relatively higher ASR@5 scores for Dutch and Spanish inputs across all judges. This strong performance can be attributed to the scarcity of safety alignment resources for relatively low-resource languages like Dutch and Spanish which is consistent with previous research Yong et al. (2023); Li et al. (2024) indicating that languages with limited alignment data typically demonstrate weaker safety measures.
@@ 3.7 Failure Analysis
By analyzing the attack success rates across categories in Figure 4, we observe distinct patterns of vulnerability when applying LogiBreak to various LLMs. Specifically, the Fraud/Deception and Privacy categories consistently exhibit high penetration rates, exceeding 80% across all models and evaluation setups, which indicates that these areas are particularly susceptible to logical adversarial prompts. In contrast, the Sexual/Adult Content and Expert Advice categories demonstrate greater resilience, with success rates reliably falling below the 80% threshold.
@@
Furthermore, the variation in attack success rates appears correlated with the perceived severity of harm associated with each category. Categories that have attracted greater public concern—such as explicit content or risky advice—tend to show stronger alignment, likely due to their more prominent representation in alignment datasets. This trend aligns with findings from Chen et al. (2024) and points to an imbalanced safety landscape, where vulnerabilities are not uniformly addressed across harm domains.
@@ 3.8 Ablation Study
In our ablation study, we further examined the unique contribution of logical translation by replacing $x_{\text{logic}}$ with the original harmful request $x_{\text{harmful}}$, while preserving all other elements of the prompt structure. As shown in Table 5, this variant reduces the ASR@5 score to a level nearly identical to that of the raw baseline reported in Table 2. This finding confirms that the jailbreak effectiveness of LogiBreak stems primarily from the logical reformulation of the input, rather than from the surrounding prompt scaffolding.
We also evaluated the impact of the prepended contextual grounding phrase $x_{\text{context}}$. When this component is removed, the attack success rate declines by approximately 8 percent on average across all evaluations, as documented in Table 5. Taken together, these ablation results demonstrate that the core efficacy of LogiBreak is driven by the logical translation step, while the inclusion of $x_{\text{context}}$ serves to further enhance performance, providing a consistent and measurable boost to overall effectiveness.
@@ 3.9 Mitigation
$X\_{logic}$ $X\_{context}$ Llama Qwen GPT-3.5 DS-V2 GPT-4o
Rule
Judge
82 78 82 88 82
81 76 76 74 73
2 10 5 28 24
Llama
Judge
70 81 83 84 85
55 72 78 71 75
0 2 0 8 1
Table 5: Ablation study of the prepended contextual grounding phrase and the role of logical translation.
We evaluate LogiBreak against two mitigation strategies: the prompt-based Self-Reminder Xie et al. (2023) technique and LLaMA-Guard-3-8B, a fine-tuned model specifically designed for safety alignment Llama Team (2024) as a filter.
For Self-Reminder mitigation against LogiBreak, effectiveness varies significantly across models. Under the Rule_Judge, only LLaMA3-8B significantly degrades, dropping from an 82% to 17% ASR, while other models remain vulnerable. With the LLaMA_Judge, LLaMA3-8B sees the best improvement, reducing its ASR from 70% to 6%; other models gain modestly, with Qwen still highly vulnerable at 70% ASR, and DeepSeek-V3, GPT-3.5, and GPT-4o-mini staying between 40-54%. Under the GPT_Judge, only LLaMA3-8B shows substantial degradation, with its ASR dropping from 48% to 6%, while all other models maintain ASRs above 20%.
For LLaMA-Guard defense, despite being fine-tuned specifically for content safety classification, approximately 36% of LogiBreak prompts successfully bypass the guard and are incorrectly classified as safe. These “safe” inputs still achieve attack success rates exceeding 30% under Rule_Judge and 20% under LLaMA_Judge and GPT_Judge evaluation across LLMs.
@@
Rule_Judge LLaMA_Judge GPT_Judge
Logi SR GD Logi SR GD Logi SR GD
LLaMA 82 17 32 70 6 22 48 6 21
Qwen 78 81 33 81 70 28 62 44 23
DS-V3 88 90 35 84 40 25 75 25 27
GPT-3.5 82 76 35 83 49 26 66 27 25
GPT-4o 82 74 34 85 54 29 63 26 27
Table 6: Performance of prompt-based Self-Reminder (SR) and finetune-based LLaMA-Guard (GD) mitigations against LogiBreak (in %).
@@ 4.1 First order logic in NLP
Our implementation of LogiBreak builds upon the established field of natural language to first-order logic (NL-FOL) translation, which has a significant role in logic-based NLP applications. In Recognizing Textual Entailment, several works have leveraged FOL to model natural language entailment Bos and Markert (2005); Lee et al. (2025). Similarly, for natural logic inference, FOL has been employed to derive commonsense conclusions Angeli and Manning (2014). Moreover, in the context of knowledge graphs, FOL enables complex query answering through embedding-based neuralization of logical operators Xia et al. (2025); Guu et al. (2015); Hamilton et al. (2018).
Early efforts in NL-to-FOL translation primarily relied on grammar-based techniques Purdy (1991); Angeli and Manning (2014); MacCartney and Manning (2014) or neural networks Singh et al. (2020); Lu et al. (2022). With the rapid progress of LLMs, however, researchers have increasingly explored LLM-based solutions. For instance, LogicLLaMA, a fine-tuned Llama variant introduced in Xiong et al. (2024), is designed specifically for NL-to-FOL conversion. The NL2FOL framework uses LLMs to convert natural language into FOL, enabling tasks like logical fallacy detection through satisfiability checking Lalwani et al. (2024). Despite these advances, the safety of FOL for LLMs remains largely unexplored, and the success of LogiBreak highlights the safety risks posed by FOL-expressed malicious requests.
@@ 4.2 Jailbreak Attack
As LLMs gain widespread use, they have been shown to be vulnerable to jailbreak attacks, where adversarial triggers induce harmful or restricted output. These attacks fall into two main types: white-box methods, which require access to model internals (e.g., logits, gradients) and are limited to open models Zou et al. (2023); Liu et al. (2024b); and black-box methods, which exploit only input–output behavior Yi et al. (2024). Due to the impracticality of white-box access, we focus on the black-box setting.
Within black-box settings, a major line of work involves prompt injection attacks, which introduce adversarial instructions to bypass safety constraints Shen et al. (2024); Wei et al. (2023); Liu et al. (2023); Zhou et al. (2024a); Yu et al. (2024). However, this is unstable and reactive as its effectiveness is easily undermined by simple model or system prompt updates, and the insufficient modifications it makes are often insufficient to bypass robust safety mechanisms. Recent works explore language-based vulnerabilities, showing that LLMs may have reduced safety in low-resource languages Yong et al. (2023); Deng et al. (2023). While this is a promising direction, it is limited by the number of languages available and cannot entirely prevent malicious content from appearing in alignment data. An alternative approach, Cipher-based methods (e.g., Base_64, Caesar) Yuan et al. (2024); Yong et al. (2023), appears to be an effective way to transform malicious requests. Nevertheless, due to the capability limitation of different models, it’s difficult to ensure a reliable encoding and decoding process, ultimately compromising semantic consistency.
In general, existing black-box jailbreak methods often fail in transforming malicious requests to bypass safety alignment without sacrificing semantic consistency, which is why they perform worse than LogiBreak.
@@ 5 Conclusion
We propose LogiBreak, a novel black-box jailbreak method that transforms malicious natural language into formal logical expressions, preserving semantic intent while evading detection by alignment mechanisms. Through experiments on multiple LLMs across three languages, we demonstrate the effectiveness of our approach, achieving high success rates with minimal attempts. The success of LogiBreak highlights a critical vulnerability in current LLMs, that is the lack of semantic-level safety alignment. Our findings emphasize the need for post-training alignment that operates at the semantic level to ensure more robust model safety.
6 Limitations
Our primary limitation lies in the constrained scope of our experimental evaluation, particularly the inability to extend testing to a broader range of jailbreak datasets. Although Logibreak has demonstrated stability and effectiveness across various languages, limited resources have restricted the scale and diversity of our assessments. Moreover, due to the limitations of population diversity in our environment, and to ensure the quality and reliability of our dataset, we opted to evaluate only five languages based on our research team’s proficiency. A more comprehensive evaluation across diverse datasets and multilingual scenarios would offer stronger evidence of Logibreak’s generalizability, which we leave as an important direction for future work.
Appendix A More Experimental Results
A.1 t-SNE visualization
Figure 5 presents a t-SNE Van der Maaten and Hinton (2008) projection of LLM2Vec BehnamGhader et al. (2024) embeddings computed for three inputs: the original natural-language request, its reformulation as a logical expression, and the whole input prompt. In both English and Chinese settings, these three embedding types form clearly separable clusters; however, the clustering is markedly sharper in the English case. We suggest that English translation systematically converts multi-word phrases into camel-cased atomic tokens (e.g., "social media campaign" to "SocialMediaCampaign"), creating semantically distinct lexical units that lack direct Chinese equivalents due to the logographic nature of Chinese script. This tokenization asymmetry manifests in the embedding space as tighter intra-cluster cohesion for English transformed tokens compared to their Chinese counterparts. These findings demonstrate that LogiBreak can successfully transform the embedding distribution, enabling effective jailbreak.
Appendix B Implementation Details
B.1 Implementation of FOL-NL Translation
In Figure 9, we provided the prompt utilized in FOL-NL Translation in five languages.
B.2 Implementation of LogiBreak
@@
B.3 Implementation of Judges
In Figures 10, 11, and 12, we showcase the prompts or functions used for the Rule-based, LLaMA, and GPT judges in three different languages.
@@ B.3 Implementation of Judges
Table 7: Categories and corresponding Abbreviations.
Appendix C Baselines
  • Casaer is a cipher-based jailbreak technique proposed in Yuan et al. (2024), which involves training the target LLM on the principles of encoding and decoding, and subsequently employing encoded malicious inputs to attempt to bypass its safety alignment mechanisms.
Figure 6: LogiBreak Prompt in English, Chinese and Dutch.
Appendix D Categories and Abbreviations in the Dataset
The request categories and their corresponding abbreviations in our dataset are presented in Table 7, following the taxonomy proposed by Chao et al. (2024).
Appendix E Instructions for Logical and Language Translation Checks
We provide in Figure 7 and Figure 8 examples of the procedures and guidelines given to volunteers for verifying the logical expression translations and the linguistic translations of malicious requests in the LogiBreak dataset, respectively. These checks aim to ensure semantic fidelity, logical correctness, and linguistic naturalness across all five target languages. All volunteers are paid \$20 per hour, which exceeds the local minimum wage.
Figure 7: Translation Accuracy Check Questionnaire.
Figure 8: Logical Translation Accuracy Check Questionnaire.
Figure 9: Prompt of FOL-NL translation for five languages